# custom_references.bib

### v1

@@ 2 Proof Overview
We give a brief overview of our strategy for our new lower bound and upper bounds for $k$-round coin flipping protocols. In Section 2.1, we sketch our proof of lower bounds for coin flipping protocols. In Section 2.2, we prove lower bounds against protocols where players can send many bits per round. In Section 2.3, we prove a useful helper theorem that shows how any family of functions can be biased by a random set of variables and a small set of variables, specific to each function. Finally, in Section 2.5, we construct new improved constant-round coin flipping protocols.
@@ 2.1 Biasing Coin Flipping Protocols
Theorem 2.1 (Two-round version of Theorem 5.1).
For any $2$-round coin flipping protocol $\pi$ over $\ell$ players where $\Pr[\pi=1]\geq 0.01$, there exists a set of bad players $B\subseteq[\ell]$ with $\left\lvert B\right\rvert=O\left(\frac{\ell}{\log\log\ell}\right)$ so that $\Pr[\pi|_{B}=1]\geq 0.99$.
@@
We will also require a simpler version of the above theorem that follows by inductively applying the KKL theorem [KKL88influence]:
@@
We will sketch a proof of Theorem 2.2 in Section 2.3. Let us see how Theorem 2.1 follows from Theorem 2.2:
@@
For $\alpha\in\{0,1\}^{\ell}$, let $\pi_{\alpha}:\{0,1\}^{\ell}\to\{0,1\}$ be the induced second round protocol when the players output $\alpha$ in the first round. Since $\Pr[\pi=1]\geq 0.01$, by a reverse Markov argument (see 3.3), we have that $\Pr[\pi_{\alpha}=1]\geq 0.001$ for at least $0.009$ fraction of $\alpha$. Let $\mathcal{F}$ be the family of functions consisting of these functions $\pi_{\alpha}$. We apply Theorem 2.2 to find a common set of bad players $B_{R}$ such that for 99% of $\pi_{\alpha}\in\mathcal{F}$, there exists a set of bad players $g(\alpha)$ such that $B_{R}\cup g(\alpha)$ can $0.999$-bias $\pi_{\alpha}$ towards $1$. Moreover, $\left\lvert B_{R}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right)$ and $\left\lvert g(\alpha)\right\rvert=h=(\log\ell)^{0.001}$.
For each $\alpha$ for which there does not exist such a set of bad players that along with $B_{R}$ can corrupt $\pi_{\alpha}$, we let $g(\alpha)=\bot$. Then, we know that $\Pr[g\neq\bot]\geq 0.008$. We use Theorem 2.3 to find a set of bad players $B_{temp}^{(1)}$ with $\left\lvert B_{temp}^{(1)}\right\rvert\leq O\left(\frac{\ell}{\log\ell}\right)$ to bias the first round protocol so that $\Pr[g|_{B_{temp}^{(1)}}\neq\bot]\geq 0.999$. We add all players from $B_{temp}^{(1)}$ to $B_{I}$.
We view $g=(g_{1},\dots,g_{h})$ where each $g_{i}:\{0,1\}^{\ell}\to[\ell]\cup\bot$ where $g_{i}$ outputs the $i$-th largest element from $g(\alpha)$ and $g_{i}(\alpha)=\bot$ if and only if $g(\alpha)=\bot$. We will repeatedly use Theorem 2.3 to find bad players that can bias each of $g_{1},\dots,g_{h}$ so that over most inputs, each of their images lies in some set of size $c=\frac{\ell}{(\log\ell)^{1000}}$. Towards this goal, we maintain sets $C_{1},\dots,C_{h}$ that will satisfy the following invariant (initially each of these sets equals $[\ell]$):
()(*)( ∗ )   For every $\alpha$, if there exists $i\in[h]$ such that $g_{i}(\alpha)\not\in C_{i}$, then $g(\alpha)=\bot$. (Equivalently, if $g(\alpha)\neq\bot$, then $g_{i}(\alpha)\in C_{i}$ for all $i\in[h].$)
We now find bad players to bias each of these $g_{i}$ and while doing so, maintain that at the beginning of each iteration of the loop, $\Pr[g|_{B_{I}}\neq\bot]\geq 0.999$. Formally, we proceed as follows:
@@
      Let $\mathbf{X}_{i}$ be a random subset of $C_{i}$ with $\left\lvert\mathbf{X}_{i}\right\rvert=\left\lvert C_{i}\right\rvert/2$. Since $\Pr[g_{i}|_{B_{I}}\neq\bot]\geq 0.999$, by our invariant property, $\Pr[{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{g_{i}|_{B_{I}}}}\in C_{i}]\geq 0.999$. Hence, $\operatorname*{\mathbb{E}}[\Pr[g_{i}|_{B_{I}}\in\mathbf{X}_{i}]]\geq 0.49$. In particular, there exists $C^{\prime}_{i}\subset C_{i}$ with $\left\lvert C^{\prime}_{i}\right\rvert=\left\lvert C^{\prime}_{i}\right\rvert/2$ such that $\Pr[g_{i}|_{B_{I}}\in C^{\prime}_{i}]\geq 0.49$. We set $C_{i}$ to equal $C^{\prime}_{i}$.
@@
      To maintain our invariant, we can now only guarantee that $\Pr[g|_{B_{I}}\neq\bot]\geq 0.49$. To increase this, we apply Theorem 2.3 to the function $g|_{B_{I}}$ to find a set of bad players $B_{temp}^{(2)}$ with $\left\lvert B_{temp}^{(2)}\right\rvert\leq O\left(\frac{\ell}{\log\ell}\right)$ so that $\Pr[g|_{B_{I}\cup B_{temp}^{(2)}}\neq\bot]\geq 0.999$. We add all players from $B_{temp}^{(2)}$ to $B_{I}$.
Finally, we let $B_{H}=\bigcup_{i=1}^{h}C_{i}$. Our final set of bad players will be $B_{R}\cup B_{H}\cup B_{I}$. We see that $\Pr[g|_{B_{I}}\neq\bot]\geq 0.999$. Whenever $g|_{B_{I}}(\alpha)\neq\bot$, $\Pr[\pi_{\alpha}|_{B_{R}\cup B_{H}}=1]\geq 0.999$. Hence, $\Pr[\pi_{B_{R}\cup B_{H}\cup B_{I}}=1]\geq 0.999\cdot 0.999\geq 0.99$ as desired.
We finally bound the number of bad players we control. We know that $\left\lvert B_{R}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right).$ We also have that $\left\lvert B_{H}\right\rvert\leq c\cdot h=\frac{\ell}{(\log\ell)^{1000}}\cdot(\log\ell)^{0.001}<\frac{\ell}{(\log\ell)^{999}}.$ Since we decrease the size of each $C_{i}$ by a factor of $2$ each time we execute the loop, and we stop once every $i\in[h]$ has $|C_{i}|\leq c$, the total number of iterations of the loop is $h\cdot\log(\ell/c)=(\log\ell)^{0.001}\cdot(1000\log\log\ell)\leq(\log\ell)^{0.002}$. Since each time through the loop we add $O\left(\frac{\ell}{\log\ell}\right)$ players to $B_{I}$, we bound the size of $B_{I}$ as $\left\lvert B_{I}\right\rvert\leq O\left(\frac{\ell}{\log\ell}+(\log\ell)^{0.002}\cdot\frac{\ell}{\log\ell}\right)\leq\frac{\ell}{(\log\ell)^{0.99}}.$ Hence, the total number of bad players we control is $\left\lvert B_{R}\right\rvert+\left\lvert B_{H}\right\rvert+\left\lvert B_{I}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right)$. ∎
We briefly mention that the above ideas more or less work to extend our two-round lower bound to $k$ rounds; the main difference is that we inductively use our bounds to bias $(k-1)$-round protocols in place of the second use case of the KKL theorem above (see Item 2(a)iii in the “Formal description of algorithm” in Section 5).
@@ 2.2 Biasing Coin Flipping Protocols with Longer Messages
We essentially follow the same proof strategy as in the $k$-round version of Theorem 2.1. To do this, we treat each round $i$ of the protocol as being over $r_{i}\cdot\ell$ bits and whenever our lower bound asks us to corrupt a bit, we corrupt the corresponding player that controls that bit. For instance, in the one-round version of this theorem, we treat the input as being over $\ell\cdot r_{1}$ bits. Then, we use our lower bounds to find a set of $\frac{\ell\cdot r_{1}}{\log(\ell\cdot r_{1})}$ bits such that if bad players control them, then they can bias the protocol. So, we let $B$ be the set of all players that control each such bit. It must be the case that $\left\lvert B\right\rvert\leq\frac{\ell\cdot r_{1}}{\log(\ell\cdot r_{1})}$. Since $\left\lvert r_{1}\right\rvert\leq(\log\ell)^{0.99}$, we end up controlling $o(\ell)$ players overall to corrupt $\pi$. This idea, with the right setting of parameters, generalizes to $k$-round protocols. ∎
@@ 2.3 Biasing a Family of Functions
[…]
We now focus on proving Lemma 2.5. Towards proving this, we will need the following result regarding influence of boolean functions that was proven in [RSZ02leader] by slightly building up on the result of [KKL88influence]:
@@
Fix $1\leq h\leq\ell^{0.99}$. Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $0.01\leq\Pr[f=1]\leq 0.99$. If $I_{f}(i)\leq\frac{1}{h}$ for all $i\in[\ell]$, then $\sum_{i=1}^{\ell}I_{f}(i)\geq\frac{\log(h)}{2000}$.
We use Lemma 2.6 to establish Lemma 2.5 by analyzing a semi-random process that was first studied in [RSZ02leader]. Our analysis closely follows that of [RSZ02leader], and only differs in the final conclusion.
@@
We say that the above process succeeds if at the end, $\Pr[f|_{B_{R}\cup B_{H}}=1]\geq 0.99$. We will establish the following:
@@
[…]
@@
Proof sketch of 2.7.
For $j\in[r]$, let $v_{j}$ represent the variable chosen at step $j$ of the semi-random process. Let $\mathbf{X}_{j}$ equal the influence of $v_{j}$ on $f|_{\{v_{1},\dots,v_{j-1}\}}$ if the process hasn’t stopped before step $j$ and let $\mathbf{X}_{j}$ equal $1$ otherwise. Let $j\in[r]$ be a step such that $\Pr[f|_{\{v_{1},\dots,v_{j-1}\}}=1]\leq 0.99$. Then, by Lemma 2.6, in step $j$, either there exists a variable with influence $\frac{2}{h}$ or else a random variable will have expected influence $\geq\frac{\log(h)}{2000\ell}$. Since $\frac{1}{h}\geq\frac{\log(h)}{2000\ell}$ (recall that we assumed $h\leq\ell^{0.99}$), we always have that $\operatorname*{\mathbb{E}}[\mathbf{X}_{j}|\mathbf{X}_{1},\dots,\mathbf{X}_{j-1}]\geq\frac{\log(h)}{2000\ell}$.
@@
We then apply Azuma’s inequality (see Lemma 4.3) to infer that $\mathbf{Z}_{r}\geq 2$ with probability at least $0.999$. Since $\mathbf{Z}_{r}$ represents the sum total of ‘contributions’ of influences of variables towards $1$, whenever $\mathbf{Z}_{r}\geq 2$ it must be that $f$ has been ‘pushed’ towards $1$ by a ‘total amount’ of $1$. In all such cases, $\Pr[f|_{B_{R}\cup B_{H}}=1]\geq 0.999$ and the process must succeed. Since this happens with probability at least $0.999$, the process succeeds with at least this probability. ∎
This concludes the proof sketch of Lemma 2.5. ∎
2.4 Comparison to Proof Strategy of [RSZ02leader]
Since our results are inspired by the work of [RSZ02leader] and our proofs build on their arguments, we here briefly sketch their strategy for 2 rounds and how we improve upon it.
Let =$\mathcal{F}=\{\pi_{\alpha}\}$ where $\alpha\in\{0,1\}^{\ell}$ and let $\pi_{\alpha}$ be the induced second round protocol when in the first round the players output $\alpha$. [RSZ02leader] proceeds by finding a set of bad players in the second round to bias the functions in $\mathcal{F}$. To do this, as hinted at earlier, [RSZ02leader] analyzes the same semi-random process as in Lemma 2.5. At the end of the process, they also obtain a common random set $B_{R}$ and a specialized heavy set $B_{H}=B_{H}(f)$ such that for most functions $f\in\mathcal{F}$, $B_{R}\cup B_{H}(f)$ can bias the function $f$. They then pick another random set $B_{R^{\prime}}$ and bound the probability that $B_{H}(f)\subset B_{R^{\prime}}$ for fixed $f$, and argue that in expectation, $\delta\approx\left(\frac{r}{\ell}\right)^{2^{\ell/r}}$ fraction of functions $f\in\mathcal{F}$ will be such that $B_{H}(f)\subset B_{R^{\prime}}$ where $\left\lvert B_{R^{\prime}}\right\rvert=r$. They let $B_{R}\cup B_{R^{\prime}}$ be the set of bad players that will bias the second round protocols.
Now, let $S\subset\{0,1\}^{\ell}$ be the set of $\alpha\in\{0,1\}^{\ell}$ such that $\pi_{\alpha}$ can be biased by $B_{R}\cup B_{R^{\prime}}$, and let $g:\{0,1\}^{\ell}\to\{0,1\}$ be such that $g(\alpha)=1\iff\alpha\in S$. We know from above that $\Pr[g=1]\geq\delta$. Now, the goal for the first round is to find bad players to bias $g$ towards $1$ so that with high probability, the resulting second round protocol can be biased using $B_{R}\cup B_{R^{\prime}}$. To do this, [RSZ02leader] use the KKL theorem [KKL88influence] which lets them do this by finding $O(\ell/(\delta\log\ell))$ bad players. To minimize the total number of bad players, they end up using $O(\ell/\log^{(3)}\ell)$ bad players. For multiple rounds, their result follows by an inductive argument, showing that $O\left(\ell/\log^{(2k-1)}\ell\right)$ bad players suffice to bias $k$ round protocols.

### v2

@@ 2 Proof Overview
In this section, we give a brief overview of our main proofs. In Section 2.1, we sketch our proof of lower bounds for coin flipping protocols. In Section 2.2, we prove lower bounds against protocols where players can send many bits per round. In Section 2.3, we show how to obtain leader election lower bounds. In Section 2.4, we prove a useful helper theorem that shows how any family of functions can be biased by a random set of variables and a small set of variables specific to each function. Then, in Section 2.5, we construct new improved constant-round coin flipping protocols. We construct one-round random selection protocols in Section 2.6. Lastly, we obtain one-round random selection lower bounds in Section 2.7.
@@ 2.1 Biasing Coin Flipping Protocols
Theorem 2.1 (Theorem 5.1, specialized).
For any two-round coin flipping protocol $\pi$ over $\ell$ players where $\Pr[\pi=1]\geq 0.01$, there exists $B\subseteq[\ell]$ with $\left\lvert B\right\rvert=O\left(\frac{\ell}{\log\log\ell}\right)$ so that $\Pr[\pi|_{B}=1]\geq 0.99$.
@@
We will also require a simpler version of the above theorem that follows from the KKL theorem [KKL88]:
@@
We will sketch a proof of Theorem 2.2 in Section 2.4. Let us see how Theorem 2.1 follows from this.
@@
For $\alpha\in\{0,1\}^{\ell}$, let $\pi_{\alpha}:\{0,1\}^{\ell}\to\{0,1\}$ be the induced second round protocol when the players output $\alpha$ in the first round. Since $\Pr[\pi=1]\geq 0.01$, by a reverse Markov argument (see 3.4), we have that $\Pr[\pi_{\alpha}=1]\geq 0.001$ for at least $0.009$ fraction of $\alpha$. Let $\mathcal{F}$ be the family of functions consisting of these functions $\pi_{\alpha}$. We apply Theorem 2.2 to find a common set of bad players $B_{R}$ such that for 99% of $\pi_{\alpha}\in\mathcal{F}$, there exists a set of bad players $g(\alpha)$ such that $B_{R}\cup g(\alpha)$ can $0.999$-bias $\pi_{\alpha}$ towards $1$. Moreover, $\left\lvert B_{R}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right)$ and $\left\lvert g(\alpha)\right\rvert=h=(\log\ell)^{0.001}$. For each $\alpha$ for which there does not exist such a set of bad players that along with $B_{R}$ can corrupt $\pi_{\alpha}$, we let $g(\alpha)=\bot$. Then, we know that $\Pr[g\neq\bot]\geq 0.008$. We use Theorem 2.3 to find a set of bad players $B_{temp}^{(1)}$ with $\left\lvert B_{temp}^{(1)}\right\rvert\leq O\left(\frac{\ell}{\log\ell}\right)$ to bias the first round protocol so that $\Pr[g|_{B_{temp}^{(1)}}\neq\bot]\geq 0.999$. We add all the players from $B_{temp}^{(1)}$ to $B_{I}$. We view $g$ as $(g_{1},\dots,g_{h})$ where each $g_{i}:\{0,1\}^{\ell}\to[\ell]\cup\{\bot\}$ where $g_{i}$ outputs the $i$-th largest element from $g(\alpha)$ and $g_{i}(\alpha)=\bot$ if and only if $g(\alpha)=\bot$. We will repeatedly use Theorem 2.3 to find bad players that can bias each of $g_{1},\dots,g_{h}$ so that over most inputs, each of their images lies in some set of size $c=\frac{\ell}{(\log\ell)^{1000}}$. Towards this goal, we maintain sets $C_{1},\dots,C_{h}$ that will satisfy the following invariant (initially each of these sets equals $\ell$]):
()(*)   For every $\alpha$, if there exists $i\in[h]$ such that $g_{i}(\alpha)\not\in C_{i}$, then $g(\alpha)=\bot$. (Equivalently, if $g(\alpha)\neq\bot$, then $g_{i}(\alpha)\in C_{i}$ for all $i\in[h].$)
We now find bad players to bias each of these $g_{i}$ and, while doing so, maintain that at the beginning of each iteration of the loop, $\Pr[g|_{B_{I}}\neq\bot]\geq 0.999$. Formally, we proceed as follows:
@@
      Let $\mathbf{X}_{i}$ be a random subset of $C_{i}$ with $\left\lvert\mathbf{X}_{i}\right\rvert=\left\lvert C_{i}\right\rvert/2$. Since $\Pr[g_{i}|_{B_{I}}\neq\bot]\geq 0.999$, by our invariant property, $\Pr[g_{i}|_{B_{I}}\in C_{i}]\geq 0.999$. Hence, $\operatorname*{\mathbb{E}}[\Pr[g_{i}|_{B_{I}}\in\mathbf{X}_{i}]]\geq 0.49$. In particular, there exists $C^{\prime}_{i}\subset C_{i}$ with $\left\lvert C^{\prime}_{i}\right\rvert=\left\lvert C_{i}\right\rvert/2$ such that $\Pr[g_{i}|_{B_{I}}\in C^{\prime}_{i}]\geq 0.49$. We set $C_{i}:=C^{\prime}_{i}$.
@@
      Now, we can only guarantee that $\Pr[g|_{B_{I}}\neq\bot]\geq 0.49$. To maintain our invariant, we apply Theorem 2.3 to the function $g|_{B_{I}}$ to find a set of bad players $B_{temp}^{(2)}$ with $\left\lvert B_{temp}^{(2)}\right\rvert\leq O\left(\frac{\ell}{\log\ell}\right)$ so that $\Pr[g|_{B_{I}\cup B_{temp}^{(2)}}\neq\bot]\geq 0.999$. We add all players from $B_{temp}^{(2)}$ to $B_{I}$.
We let $B_{H}=\bigcup_{i=1}^{h}C_{i}$. Our final set of bad players will be $B_{R}\cup B_{H}\cup B_{I}$. We see that $\Pr[g|_{B_{I}}\neq\bot]\geq 0.999$. Whenever $g|_{B_{I}}(\alpha)\neq\bot$, $\Pr[\pi_{\alpha}|_{B_{R}\cup B_{H}}=1]\geq 0.999$. Hence, $\Pr[\pi_{B_{R}\cup B_{H}\cup B_{I}}=1]\geq 0.999\cdot 0.999\geq 0.99$ as desired.
We finally bound the number of bad players that we control. We know that $\left\lvert B_{R}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right).$ We also have that $\left\lvert B_{H}\right\rvert\leq c\cdot h=\frac{\ell}{(\log\ell)^{1000}}\cdot(\log\ell)^{0.001}<\frac{\ell}{(\log\ell)^{999}}.$ Since we decrease the size of each $C_{i}$ by a factor of $2$ each time we execute the loop, and we stop once every $i\in[h]$ has $|C_{i}|\leq c$, the total number of iterations of the loop is $h\cdot\log(\ell/c)=(\log\ell)^{0.001}\cdot(1000\log\log\ell)\leq(\log\ell)^{0.002}$. Since each time through the loop we add $O\left(\frac{\ell}{\log\ell}\right)$ players to $B_{I}$, we bound the size of $B_{I}$ as $\left\lvert B_{I}\right\rvert\leq O\left(\frac{\ell}{\log\ell}+(\log\ell)^{0.002}\cdot\frac{\ell}{\log\ell}\right)\leq\frac{\ell}{(\log\ell)^{0.99}}.$ Hence, the total number of bad players we control is $\left\lvert B_{R}\right\rvert+\left\lvert B_{H}\right\rvert+\left\lvert B_{I}\right\rvert\leq O\left(\frac{\ell}{\log\log\ell}\right)$. ∎
The above ideas more or less work to extend our two-round lower bound to $k$ rounds; the main difference is that we inductively use our bounds to bias (k1)(k-1)-round protocols in place of the second use case of the KKL theorem above (see Item 2(a)iii in the “Formal description of algorithm” in Section 5).
@@ 2.2 Biasing Coin Flipping Protocols with Longer Messages
We essentially follow the same proof strategy as in the $k$-round version of Theorem 2.1. To do this, we treat each round $i$ of the protocol as being over $r_{i}\cdot\ell$ bits and whenever our lower bound asks us to corrupt a bit, we corrupt the corresponding player that controls that bit. For instance, in the one-round version of this theorem, we treat the input as being over $\ell\cdot r_{1}$ bits. Then, we use our lower bounds to find a set of $\frac{\ell\cdot r_{1}}{\log(\ell\cdot r_{1})}$ bits such that if bad players control them, then they can bias the protocol. So, we let $B$ be the set of all players that control each such bit. It must be the case that $\left\lvert B\right\rvert\leq\frac{\ell\cdot r_{1}}{\log(\ell\cdot r_{1})}$. Since $\left\lvert r_{1}\right\rvert\leq(\log\ell)^{0.99}$, we end up controlling $o(\ell)$ players overall to corrupt $\pi$. This generalizes to $k$-round protocols. ∎
2.3 Biasing Leader Election Protocols
Here we give a sketch of how to obtain our leader election lower bounds using coin flipping impossibility results. We will show this for one round and the idea for $k$ rounds will follow similarly.
Theorem 2.5 (One-round version of Corollary 5.3).
For all one-round leader election protocols $\pi$, there exist $b=O\left(\frac{\ell\log(\log(\ell))}{\log(\ell)}\right)$ bad players that can corrupt $\pi$ to elect a bad leader with probability $0.99$.
Proof.
We will repeatedly apply the one round coin flipping impossibility result given by Theorem 2.3. Let $B$ be our collection of bad players, initialized to $\varnothing$, let $H_{0}=[\ell]$, and let $t=\log(\log(\ell))$. For $i\in[t]$, we will maintain that the players in $B$ can corrupt $\pi$ to output from $H_{i}$ with probability $\geq 0.99$. Using this guarantee, at step $i\in[t]$ we find $H_{i}\subset H_{i-1}$ with $\left\lvert H_{i}\right\rvert=\left\lvert H_{i-1}\right\rvert/2$ such that the players in $B$ can corrupt $\pi$ to output from $H_{i-1}$ with probability at least $0.49$. We then consider a one-round coin flipping protocol that outputs $1$ if $B$ succeeds in corrupting $\pi$ to output from $H_{i}$ and outputs 0 otherwise. Using Theorem 2.3, we corrupt this coin flipping protocol using $O\left(\frac{\ell}{\log(\ell)}\right)$ bad players, add them to $B$, and meet the guarantee for round $i$ that we output from $H_{i}$ with probability $0.99$. The final size of $B$ will be $O\left(\frac{\ell}{\log(\ell)}\cdot t\right)=O\left(\frac{\ell\log(\log(\ell))}{\log(\ell)}\right)$, and the final size of $H_{t}$ will be $\ell/2^{t}=\frac{\ell}{\log(\ell)}$. We add $H$ to $B$ to infer the claim. ∎
@@ 2.4 Biasing a Family of Functions
We set $h=(\log\ell)^{0.001}$. Let $G=(U,V)$ be a bipartite graph where the left part $U$ equals $\mathcal{F}$, and the elements of the right part $V$ are all size $\frac{\ell}{\log(h)}$ subsets of $\ell$]. We add an edge between $f\in U$ and a set $R\in V$ if there exists $H\subset[\ell]$ with $\left\lvert H\right\rvert\leq h$ such that $R\cup H$ can $0.99$-bias $f$ towards $1$. By Lemma 2.6, the degree of each $f\in U$ is at least $0.99\cdot\left\lvert V\right\rvert$. Therefore, there exists $B_{R}\in V$ such that $B_{R}$ has degree at least $0.99\cdot\left\lvert U\right\rvert$. Such a $B_{R}$ satisfies the conditions of the theorem. ∎
We now focus on proving Lemma 2.6. Towards proving this, we will need the following result regarding the influence of boolean functions that was proven in [RSZ02] by slightly building up on the result of [KKL88].
@@
Fix $1\leq h\leq\ell^{0.99}$. Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $0.01\leq\Pr[f=1]\leq 0.99$. If $\operatorname{\mathbf{I}}_{i}(f)\leq\frac{1}{h}$ for all $i\in[\ell]$, then $\sum_{i=1}^{\ell}\operatorname{\mathbf{I}}_{i}(f)\geq\frac{\log(h)}{2000}$, where $\operatorname{\mathbf{I}}_{i}(f)$ is the influence of the coordinate $i$ on $f$.
We use Lemma 2.7 to establish Lemma 2.6 by analyzing a semi-random process first studied in [RSZ02]. Our analysis closely follows that of [RSZ02], and only differs in the final conclusion.
@@
We say that the above process succeeds if at the end, $\Pr[f|_{B_{R}\cup B_{H}}=1]\geq 0.99$. We claim the following:
@@
We will prove this claim later. For now we show how Lemma 2.6 follows from it. First, we observe that $\left\lvert B_{H}\right\rvert\leq h$ (since each player added to $B_{H}$ “pushes” $f$ towards $1$ by a factor of $1/h$, this can happen at most $h$ times). Next, consider a random set $\mathbf{R}\subset[\ell]$ with $\left\lvert\mathbf{R}\right\rvert$$\ell$ and a random permutation $\boldsymbol{\pi}$ of $[r]$ so that $(\mathbf{R},\boldsymbol{\pi})$ fixes an ordering of the elements of $\mathbf{R}$. Consider the modified semi-random process where an initially randomly choice of $(\mathbf{R},\boldsymbol{\pi})$ is made, and then in the random case of the process, the earliest element from $\mathbf{R}$ that is not in $B_{R}\cup B_{H}$ is chosen. At the end of the process, $B_{R}\subset\mathbf{R}$, and if the process succeeds, $B_{H}\cup\mathbf{R}$ can indeed bias $f$, as desired. By 2.8, the process succeeds with probability at least $0.999$, so for a $0.999$ fraction of choices of $(R,\pi)$, the modified process succeeds. Consequently, there exists a $\pi^{*}$ such that for a $0.999$ fraction of $(R,\pi^{*})$, the modified random process succeeds. Thus, for a $0.999$ fraction of sets $R$, there exists a set $B_{H}=B_{H}(R)$ (given by the process) such that $R\cup B_{H}$ can $0.99$ bias $f$ as desired. ∎
@@
Proof sketch of 2.8.
For $j\in[r]$, let $v_{j}$ represent the variable chosen at step $j$ of the semi-random process. Let $\mathbf{X}_{j}$ equal the influence of $v_{j}$ on $f|_{\{v_{1},\dots,v_{j-1}\}}$ if the process hasn’t stopped before step $j$, and let $\mathbf{X}_{j}$ equal $1$ otherwise. Let $j\in[r]$ be a step such that $\Pr[f|_{\{v_{1},\dots,v_{j-1}\}}=1]\leq 0.99$. Then, by Lemma 2.7, in step $j$ there either exists a variable with influence $\frac{2}{h}$ or else a randomly chosen variable will have expected influence $\geq\frac{\log(h)}{2000\ell}$. Since $\frac{1}{h}\geq\frac{\log(h)}{2000\ell}$ (recall that we assumed $h\leq\ell^{0.99}$), we always have that $\operatorname*{\mathbb{E}}[\mathbf{X}_{j}|\mathbf{X}_{1},\dots,\mathbf{X}_{j-1}]\geq\frac{\log(h)}{2000\ell}$.
@@
We apply Azuma’s inequality (see Lemma 4.3) to infer that $\mathbf{Z}_{r}\geq 2$ with probability $\geq 0.999$. Since $\mathbf{Z}_{r}$ represents the sum total of ‘contributions’ of influences towards $1$, whenever $\mathbf{Z}_{r}\geq 2$ it must be that $f$ has been ‘pushed’ towards $1$ by a total amount of $1$; so, $\Pr[f|_{B_{R}\cup B_{H}}=1]\geq 0.999$ and the process succeeds. Since this happens with probability $\geq 0.999$, the process succeeds with at least this probability. ∎